OpenStudy (anonymous):

What is the sum of a 12–term arithmetic sequence where the last term is 13 and the common difference is –10? is it 816?

OpenStudy (anonymous):

\[13 = a + 11*(-10)\] \[a = 13 + 110 = 123\] \[S_{12} = \frac{12}{2} (123+ 13)\] \[S_{12} = 6 \times 136 \]

OpenStudy (anonymous):

thank you